# Automated generation of photonic circuits for Bell tests with homodyne measurements

### v1

@@ A.1 Gaussian states
We consider circuits composed of $N$ bosonic modes. Each mode is identified with an index $i\in\{1,\dots,N\}$ and is associated with the ladder operators $\hat{a}_{i}$ and $\hat{a}_{i}^{\dagger}$ or, alternatively, to the dimensionless quadrature field operators $\hat{x}_{i}=\frac{\hat{a}_{i}^{\dagger}+\hat{a}_{i}}{2}$ and $\hat{p}_{i}=i\frac{\hat{a}_{i}^{\dagger}-\hat{a}_{i}}{2}$, which satisfy the commutation relation: $[\hat{x}_{i},\hat{p}_{i}]=i\frac{1}{2}$. By arranging the quadrature field operators in the vector: $\mathbf{\hat{q}}=(\hat{x}_{1},\hat{p}_{1},\ldots,\hat{x}_{N},\hat{p}_{N})$, we have the commutation relation
@@
$\Omega=\bigoplus_{i=1}^{N}\omega,\quad\omega=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}.$ (15)
@@
where $\text{d}^{2}\beta=\text{d}\text{Re}(\beta)\,\text{d}\text{Im}(\beta)$. A $N$ bosonic modes quantum state can be expressed in terms of its Wigner function following
$\hat{\rho}=\pi^{N}\int d^{2N}\tilde{{\bm{\alpha}}}\,W(\tilde{{\bm{\alpha}}})\,\bigotimes_{i=1}^{N}\delta(\hat{a}_{i}-\alpha_{i}).$ (18)
@@
$W_{{\bm{\mu}},\Sigma}(\tilde{{\bm{\alpha}}})$ $=\frac{\exp\left[-\frac{1}{2}(\tilde{}{\bm{\alpha}}-{\bm{\mu}})^{T}\Sigma^{-1}(\tilde{}{\bm{\alpha}}-{\bm{\mu}})\right]}{(2\pi)^{n}\sqrt{\det\Sigma}},$ (19)
@@
$\mu_{i}$ $=\langle{\hat{q}_{i}}\rangle,$ (20)
$\Sigma_{i,j}$ $=\frac{1}{2}\langle{\hat{q_{i}}\hat{q_{j}}+\hat{q_{j}}\hat{q_{i}}}\rangle-\mu_{i}\mu_{j}.$ (21)
Note that, due to the commutation relations in Eq. 14, an arbitrary real symmetric matrix $\Sigma$ is a valid covariance matrix if and only if the inequality
@@ A.2 Gaussian operations
In Table 1 we report the Gaussian operations, also named gates, that are considered in this manuscript, in terms of both the unitary transformation and the corresponding vector $\bm{d}$ and symplectic matrix $M$. We adopt the notation for which the square brackets $[i,j]$ indicate that the gate acts non-trivially on modes $i$ and $j$, with $i<j$. Note that we restrict the parameters of each operation to the real domain, as any Gaussian operation with a complex parameter can be achieved with a combination of Gaussian operations with real parameters and phase shifters.
@@
Phase Shifter $\hat{R}(\theta)[i]=e^{-i\theta\hat{a}^{\dagger}_{i}\hat{a}_{i}}$ $M=\left(\begin{array}[]{c|c|c}\mathbb{I}&0&0\\ \hline\cr 0&\begin{matrix}\cos(\theta)&\sin(\theta)\\ -\sin(\theta)&\cos(\theta)\\ \end{matrix}&0\\ \hline\cr 0&0&\mathbb{I}\end{array}\right)$
Single Mode Squeezer $\hat{S}_{1}(r)[i]=e^{r[\hat{a}_{i}^{2}-(\hat{a}_{i}^{\dagger})^{2}]/2}$ $M=\left(\begin{array}[]{c|c|c}\mathbb{I}&0&0\\ \hline\cr 0&\begin{matrix}e^{-r}&0\\ 0&e^{r}\\ \end{matrix}&0\\ \hline\cr 0&0&\mathbb{I}\end{array}\right)$
Beam Splitter $\hat{B}(\theta)[i,j]=e^{\theta(\hat{a}_{i}^{\dagger}\hat{a}_{j}-\hat{a}_{i}\hat{a}_{j}^{\dagger})}$ $M=\left(\begin{array}[]{c|c|c|c|c}\mathbb{I}&0&0&0&0\\ \hline\cr 0&\cos(\theta)\,\mathbb{I}_{2}&0&\sin(\theta)\,\mathbb{I}_{2}&0\\ \hline\cr 0&0&\mathbb{I}&0&0\\ \hline\cr 0&-\sin(\theta)\,\mathbb{I}_{2}&0&\cos(\theta)\,\mathbb{I}_{2}&0\\ \hline\cr 0&0&0&0&\mathbb{I}\end{array}\right)$
Two Modes Squeezer $\hat{S}_{2}(r)[i,j]=e^{r(\hat{a}_{i}\hat{a}_{j}-\hat{a}_{i}^{\dagger}\hat{a}_{j}^{\dagger})}$ $M=\left(\begin{array}[]{c|c|c|c|c}\mathbb{I}&0&0&0&0\\ \hline\cr 0&\cosh(r)\,\mathbb{I}_{2}&0&-\sinh(r)\,\mathbb{I}_{2}&0\\ \hline\cr 0&0&\mathbb{I}&0&0\\ \hline\cr 0&-\sinh(r)\,\mathbb{I}_{2}&0&\cosh(r)\,\mathbb{I}_{2}&0\\ \hline\cr 0&0&0&0&\mathbb{I}\end{array}\right)$
Table 1: The Gaussian operations considered in this manuscript, when acting on an $N$-mode Gaussian state. Here $\mathbb{I}_{n}$ is the $n\times n$ identity matrix and $\sigma_{z}$ is the third Pauli matrix. Note that the non-trivial $2\times 2$ blocks in the symplectic matrices correspond to the modes $i$ and $j$, i.e. to rows/columns pairs $2i-1,2i$ and $2j-1,2j$. For all the operations considered, $\bm{d}=0$.
@@ A.3 Photon heralding
${\bm{\mu}}_{\circ i}$ $=\mathrm{TR}_{i}[(\Sigma^{-1}+F)^{-1}\Sigma^{-1}{\bm{\mu}}]$
$\Sigma_{\circ i}$ $=\mathrm{TR}_{i}[(\Sigma^{-1}+F)^{-1}],$ (27)
where the $2N\times 2N$ matrix $F$ is null everywhere except for the elements at position $12i-$ 1 and $2i$, which we write
@@
$p_{\circ i}=\frac{2}{2-\eta}\sqrt{\frac{(\det\Sigma)^{-1}}{\det(\Sigma^{-1}+F)}}\,e^{-\frac{1}{2}{\bm{\mu}}^{T}\left[\Sigma^{-1}-\Sigma^{-1}(\Sigma^{-1}+F)^{-1}\Sigma^{-1}\right]{\bm{\mu}}}.$ (32)
@@
$\hat{\rho}_{\bullet i}=\frac{1}{p_{\bullet i}}\hat{\rho}_{\neg i}-\frac{p_{\circ i}}{p_{\bullet i}}\hat{\rho}_{\circ i},$ (33)
where the heralding probability is $p_{\bullet i}=1-p_{\circ i}$, with $p_{\circ i}$ defined in Eq. 32.
@@ A.4 CHSH score from homodyne measurements
where in the first equality we express the Wigner function of the heralded state as a linear combination of Gaussian Wigner functions, $W^{k}_{12}$, according to Eq. 10, while in the second equality we compute the marginal of a Gaussian distribution. In particular, the $2\times 2$ matrix $\sigma_{k}$ is obtained from the covariance matrix of the $k$-th Guassian state by removing the second and fourth row and column, while the 2-dimensional vector ${\bm{\mu}}_{k}$ is obtained from the first and third entry of the displacement vector of the $k$-th Gaussian state. Finally, we have that $\tilde{{\bm{\alpha}}}^{T}=(x_{1},x_{2})$.
Although the matrices $\sigma_{k}$ are symmetric and satisfy $\sigma_{k}\geq 0$, the exponential functions in Eq. 36 are not necessarily Gaussian. This is because integrating a Gaussian Wigner function over the variables $p_{1}$ and $p_{2}$ does not always result in a Gaussian distribution. In particular, if the matrices $\sigma_{k}$ are singular, the expression becomes undefined. However, in practice, the determinants are never exactly zero, allowing us to safely use the expression in Eq. 36.
Now, let us consider a general binning function $a(x_{1}):\mathbb{R}\to\{-1,1\}$ for Alice and $b(x_{2}):\mathbb{R}\to\{-1,1\}$ for Bob. The special case of sign binning is recovered for $a(x)=b(x)=1$ for $x>0$ and $a(x)=b(x)=-1$ otherwise. Then, the generic correlator appearing in the CHSH score is expressed in terms of the joint probability $P(x_{1},x_{2})$ as
@@
By employing the result in Eq. 36, we can recast the above expression as
@@
Note that the region $R_{+}$ is given by a set of disjoint rectangular areas $R_{\square}=[u_{1},v_{1}]\times[u_{2},v_{2}]$ that depend on the specific choice of binning functions $a(x)$ and $b(x)$. Thus, for the sake of calculating Eq. 39, we are interested in the integral
@@
By partially integrating over $x_{1}$, one can show that the integral in Eq. 40 reduces to
@@
where the quantities $I_{\square}$ are obtained via Eq. 42 and read
@@
One can then employ Eq. 44, in combination with a numerical computation of Eq. 46, to compute the CHSH score for any arbitrary choice of binning functions. For the specific choice that we made in our work, i.e. sign binning, we have
@@
where $I_{k}$ and $I^{\prime}_{k}$ are obtained by numerically solving Eq. 46 with parameters $u_{1}=u_{2}=0$ and $v_{1}=v_{2}=\infty$ for $I_{k}$ and $u_{1}=u_{2}=-\infty$ and $v_{1}=v_{2}=0$ for $I^{\prime}_{k}$.
@@ B.1 Exploring photonic setups
We explore $N$-mode photonic circuits composed of $n_{\mathrm{circuit}}$ optical elements. Photonic operations are applied to the $N$-modes during a state preparation phase. The last $N-2$ modes are then heralded, while the first two modes are sent to Alice and Bob respectively.
To enhance the experimental relevance of discovered circuit and the chance to find interesting setups, we devise exploration strategies ; constraints on the initial state and on the type of the available photonic operations. The strategies are detailed in Table 2. These strategies are physically-informed, they leverage knowledge on photonic state construction and answer practical experimental requirements.
We remark that all strategies but Strategy 1 are characterized by having the initial vacuum modes initialized by active gates $S_{1}$ and $S_{2}$, followed by only passive gates. This approach has some advantages. First of all, by virtue of Euler’s decomposition in Eq. 26, this approach is non-restrictive as it still enables us to explore the whole set of circuits that can be generated by the full set of gates. Second, we avoid circuits with multiple squeezers in sequence, which not only may be experimentally challenging but can also cause numerical errors due to a compounded large squeezing parameter, thereby invalidating the results.
@@ B.2 Results
We run the random search and reinforcement learning routine over a fixed number of episodes, where each episode terminates when the circuit depth reaches its maximum predefined value, $n_{\rm circuit}$. We explored different values of $n_{\rm circuit}$ to verify what violations can be achieved both in circuits with few gates and in circuits with large depth. The photonic gates parameters are optimized at the end of each episode such that the CHSH score is maximized, with a numerical procedure based on Nelder-Mead optimization. Subsequently, we fix the CHSH score and optimize the gates parameters to maximize the total heralding probability of the state.
In Table 3 and Table 4, we report a selection of photonic circuits found by random search and reinforcement learning, characterized by their CHSH score and heralding probability. We focus on circuits with $N=4$ and $N=6$ modes, as our findings suggest that five modes offer no significant advantage over four.
In order to evaluate the various strategies without having to fine tune the RL parameters (e.g. update frequency, trajectory capacity, clip range, ADAM optimizer, and so on), we perform a random search, where the photonic operations are selected uniformly at random in each step, thereby saving us some time. In this evaluation, strategies 2 and 4 appear suboptimal (c.f. Table 3), and we thus focus on the other strategies.
@@
Figure 5: The evolution of the CHSH score with the number of completed episodes, for an agent with PPO policy implementing three different strategies to build the optical circuits (c.f. Appendix B). In both (a) and (b), the increase of the CHSH score with the number of episodes shows that the agent learns from previous interactions with the environment to produce circuits with higher CHSH violations. In particular, the agent finds circuits with CHSH score equal to 2.068 (red dashed line) in (a) and 2.072 (red dashed line) in (b). Conversely, the CHSH score remains constantly below 2.0 (red line) in (c), signifying that the agent could not learn a policy leading to CHSH violations. Samples of the circuits found with the three strategies are reported in Table 4.
@@
The circuits found by the agent in Table 4 are representative of the learning process occurring with the PPO policy. In fact, Table 4 reports circuits found by the agent towards the end of the training process, where the CHSH score associated with the circuit becomes larger. The learning process can be visualized with the plots in Fig. 5, which show the progression of the CHSH score over the training episodes. We observe that, for Strategy 1 in Fig. 5(a) and Strategy 3 in Fig. 5(b), the CHSH score improves over the training and reaches the maximum value in the later episodes. Conversely, it seems that the PPO agent struggles to learn circuit patterns leading to CHSH violations when adopting Strategy 5. This could be explained by the fact that, in Strategy 5, Alice’s and Bob’s modes are initialized in a product state that would result, if measured, in uncorrelated measurement outcomes ($\braket{A_{x}B_{y}}=0$), opposed to the correlations that arise when initializing the modes with a two-mode squeezer ($\braket{A_{x}B_{y}}>0$). This means that the starting value of the CHSH score, when the circuit is still empty, with Strategy 5 is closer to zero compared to Strategy 3, where it starts closer to $2$. Thus, it is easier for the agent in Strategy 3 to find circuit patterns leading to CHSH violations than when Strategy 5 is adopted. Still, we observe that the PPO agent in Strategy 5 can find circuits leading to a CHSH score of 2.0, i.e., a policy that generate (trivial) classical correlations between Alice and Bob is learned.
@@
Finally, training the agent both with the PPO policy and the random search allows us to benchmark the usefulness of a learning agent for circuit generation. By comparing the circuits found in Tables 3 and 4, there is no apparent advantage in using a trained agent to find good candidate circuits, compared to a random search, in this scenario. We remark, however, that this conclusion holds in the context of our RL trainings, which are limited by the computational power of a personal computer and by a shallow neural network with two hidden layers, but we expect a larger neural network to perform better than the one used here. Furthermore, we expect the performance of an unstructured search to be less appealing when trying to construct circuits in larger scenarios, e.g. involving more than two parties.
@@
Table 3: Circuits found by the agent with random search. “S.” indicates the strategy adopted, $g$ is the number of Gaussian operations, where we indicate in brackets the number of active Gaussian operations (squeezers). The notation is the one adopted in Table 1 and the gates are applied sequentially in the same order as they appear in the list. We omit the $\hat{}$ symbol in the gates list for ease of notation. The heralding of the modes $3,4,\dots,N$ is “cl.” when the state is heralded on a click in a threshold detector, and “s.p.” when attained by simulating a single photon projection; we also report the total heralding probability of the state under “Herald.”.
S. $N$ $g$ Gates Herald. CHSH
@@
S. $N$ $g$ Gates Herald. CHSH
@@
Table 4: Circuits found using reinforcement learning with the PPO policy. The same notation of Table 3 applies.
S. $N$ $g$ Gates Herald. CHSH

### v2

@@ A.1 Gaussian states
We consider circuits composed of $N$ bosonic modes. Each mode is identified with an index $l\in\{1,\dots,N\}$ and is associated with the ladder operators $\hat{a}_{l}$ and $\hat{a}_{l}^{\dagger}$ or, alternatively, to the dimensionless quadrature field operators $\hat{x}_{l}=\frac{\hat{a}_{l}^{\dagger}+\hat{a}_{l}}{2}$ and $\hat{p}_{l}=i\frac{\hat{a}_{l}^{\dagger}-\hat{a}_{l}}{2}$, which satisfy the commutation relation: $\hat{x}_{l},\hat{p}_{l}]=i\frac{1}{2}$. By arranging the quadrature field operators in the vector: $\mathbf{\hat{q}}=(\hat{x}_{1},\hat{p}_{1},\ldots,\hat{x}_{N},\hat{p}_{N})$, we have the commutation relation
@@
$\Omega=\bigoplus_{l=1}^{N}\omega,\quad\omega=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$. (15)
@@
where $\text{d}^{2}\beta=\text{d}\text{Re}(\beta)\,\text{d}\text{Im}(\beta)$. A quantum state on $N$ bosonic modes can be expressed in terms of its Wigner function following
$\hat{\rho}=\pi^{N}\int d^{2N}\tilde{{\bm{\alpha}}}\,W(\tilde{{\bm{\alpha}}})\,\bigotimes_{l=1}^{N}\delta(\hat{a}_{l}-\alpha_{l}).$ (18)
@@
[…]
@@
$\mu_{j}$ $=\langle{\hat{q}_{j}}\rangle,$ (20)
$\Sigma_{j,k}$ $=\frac{1}{2}\langle{\hat{q_{j}}\hat{q_{k}}+\hat{q_{k}}\hat{q_{j}}}\rangle-\mu_{j}\mu_{k}.$ (21)
Note that, due to the commutation relations in Eq.˜14, an arbitrary real symmetric matrix $\Sigma$ is a valid covariance matrix if and only if the inequality
@@ A.2 Gaussian operations
In Table˜1 we report the Gaussian operations, also named gates, that are considered in this manuscript, in terms of both the unitary transformation and the corresponding vector $\bm{d}$ and symplectic matrix $M$. We adopt the notation for which the square brackets $[j,k]$ indicate that the gate acts non-trivially on modes $j$ and $k$, with $j<k$. Note that we restrict the parameters of each operation to the real domain, as any Gaussian operation with a complex parameter can be achieved with a combination of Gaussian operations with real parameters and phase shifters.
@@
Phase Shifter $\hat{R}(\theta)[l]=e^{-i\theta\hat{a}^{\dagger}_{l}\hat{a}_{l}}$ $M=\left(\begin{array}[]{c|c|c}\mathbb{I}&0&0\\ \hline\cr 0&\begin{matrix}\cos(\theta)&\sin(\theta)\\ -\sin(\theta)&\cos(\theta)\\ \end{matrix}&0\\ \hline\cr 0&0&\mathbb{I}\end{array}\right)$
Single Mode Squeezer $\hat{S}_{1}(r)[l]=e^{r[\hat{a}_{l}^{2}-(\hat{a}_{l}^{\dagger})^{2}]/2}$ $M=\left(\begin{array}[]{c|c|c}\mathbb{I}&0&0\\ \hline\cr 0&\begin{matrix}e^{-r}&0\\ 0&e^{r}\\ \end{matrix}&0\\ \hline\cr 0&0&\mathbb{I}\end{array}\right)$
Beam Splitter $\hat{B}(\theta)[j,k]=e^{\theta(\hat{a}_{j}^{\dagger}\hat{a}_{k}-\hat{a}_{j}\hat{a}_{k}^{\dagger})}$ $M=\left(\begin{array}[]{c|c|c|c|c}\mathbb{I}&0&0&0&0\\ \hline\cr 0&\cos(\theta)\,\mathbb{I}_{2}&0&\sin(\theta)\,\mathbb{I}_{2}&0\\ \hline\cr 0&0&\mathbb{I}&0&0\\ \hline\cr 0&-\sin(\theta)\,\mathbb{I}_{2}&0&\cos(\theta)\,\mathbb{I}_{2}&0\\ \hline\cr 0&0&0&0&\mathbb{I}\end{array}\right)$
Two Modes Squeezer $\hat{S}_{2}(r)[j,k]=e^{r(\hat{a}_{j}\hat{a}_{k}-\hat{a}_{j}^{\dagger}\hat{a}_{k}^{\dagger})}$ $M=\left(\begin{array}[]{c|c|c|c|c}\mathbb{I}&0&0&0&0\\ \hline\cr 0&\cosh(r)\,\mathbb{I}_{2}&0&-\sinh(r)\,\mathbb{I}_{2}&0\\ \hline\cr 0&0&\mathbb{I}&0&0\\ \hline\cr 0&-\sinh(r)\,\mathbb{I}_{2}&0&\cosh(r)\,\mathbb{I}_{2}&0\\ \hline\cr 0&0&0&0&\mathbb{I}\end{array}\right)$
Table 1: The Gaussian operations considered in this manuscript, when acting on an $N$-mode Gaussian state. Here $\mathbb{I}_{n}$ is the $n\times n$ identity matrix and $\sigma_{z}$ is the third Pauli matrix. Note that the non-trivial $2\times 2$ blocks in the symplectic matrices correspond to the modes $j$ and $k$, i.e. to rows/columns pairs 2j1,2j2j-1,2j and 2k1,2k2k-1,2k. For all the operations considered, $\bm{d}=0$.
@@ A.3 Photon heralding
${\bm{\mu}}_{\circ l}$ $=\mathrm{TR}_{l}[(\Sigma^{-1}+F)^{-1}\Sigma^{-1}{\bm{\mu}}]$
$\Sigma_{\circ l}$ $=\mathrm{TR}_{l}[(\Sigma^{-1}+F)^{-1}],$ (27)
where the $2N\times 2N$ matrix $F$ is null everywhere except for the elements at position 2l12l-1 and $2l$, which we write
@@
$p_{\circ l}=\frac{2}{2-\eta}\sqrt{\frac{(\det\Sigma)^{-1}}{\det(\Sigma^{-1}+F)}}\,e^{-\frac{1}{2}{\bm{\mu}}^{T}\left[\Sigma^{-1}-\Sigma^{-1}(\Sigma^{-1}+F)^{-1}\Sigma^{-1}\right]{\bm{\mu}}}.$ (32)
@@
$\hat{\rho}_{\bullet l}=\frac{1}{p_{\bullet l}}\hat{\rho}_{\neg l}-\frac{p_{\circ l}}{p_{\bullet l}}\hat{\rho}_{\circ l},$ (33)
where the heralding probability is $p_{\bullet l}=1-p_{\circ l}$, with $p_{\circ l}$ defined in Eq.˜32.
@@ A.4 CHSH score from homodyne measurements
where in the first equality we express the Wigner function of the heralded state as a linear combination of Gaussian Wigner functions, $W^{k}_{12}$, according to Eq.˜10, while in the second equality we compute the marginal of a Gaussian distribution. In particular, the $2\times 2$ matrix $\sigma_{k}$ is obtained from the covariance matrix of the $k$-th Guassian state by removing the second and fourth row and column, while the 2-dimensional vector ${\bm{\mu}}_{k}$ is obtained from the first and third entry of the displacement vector of the $k$-th Gaussian state. Finally, we have that $\tilde{{\bm{\alpha}}}^{T}=(x_{1},x_{2})$.
Although the matrices $\sigma_{k}$ are symmetric and satisfy $\sigma_{k}\geq 0$, the exponential functions in Eq.˜36 are not necessarily Gaussian. This is because integrating a Gaussian Wigner function over the variables $p_{1}$ and $p_{2}$ does not always result in a Gaussian distribution. In particular, if the matrices $\sigma_{k}$ are singular, the expression becomes undefined. However, in practice, the determinants are never exactly zero, allowing us to safely use the expression in Eq.˜36.
Now, let us consider a general binning function $a(x_{1}):\mathbb{R}\to\{-1,1\}$ for Alice and $b(x_{2}):\mathbb{R}\to\{-1,1\}$ for Bob. The special case of sign binning is recovered for $a(x)=b(x)=1$ for $x>0$ and a(x)=b(x)=1a(x)=b(x)=-1 otherwise. Then, the generic correlator appearing in the CHSH score is expressed in terms of the joint probability $P(x_{1},x_{2})$ as
@@
By employing the result in Eq.˜36, we can recast the above expression as
@@
Note that the region $R_{+}$ is given by a set of disjoint rectangular areas $R_{\square}=[u_{1},v_{1}]\times[u_{2},v_{2}]$ that depend on the specific choice of binning functions $a(x)$ and $b(x)$. Thus, for the sake of calculating Eq.˜39, we are interested in the integral
@@
By partially integrating over $x_{1}$, one can show that the integral in Eq.˜40 reduces to
@@
where the quantities $I_{\square}$ are obtained via Eq.˜42 and read
@@
One can then employ Eq.˜44, in combination with a numerical computation of Eq.˜46, to compute the CHSH score for any arbitrary choice of binning functions. For the specific choice that we made in our work, i.e. sign binning, we have
@@
where $I_{k}$ and $I^{\prime}_{k}$ are obtained by numerically solving Eq.˜46 with parameters $u_{1}=u_{2}=0$ and $v_{1}=v_{2}=\infty$ for $I_{k}$ and $u_{1}=u_{2}=$-$\infty$ and $v_{1}=v_{2}=0$ for $I^{\prime}_{k}$.
@@ B.1 Exploring photonic setups
We explore $N$-mode photonic circuits composed of $n_{\mathrm{circuit}}$ optical elements. Photonic operations are applied to the $N$-modes during a state preparation phase. The last N2N-2 modes are then heralded, while the first two modes are sent to Alice and Bob respectively.
To enhance the experimental relevance of discovered circuit and the chance to find interesting setups, we devise exploration strategies ; constraints on the initial state and on the type of the available photonic operations. The strategies are detailed in Table˜2. These strategies are physically-informed, they leverage knowledge on photonic state construction and answer practical experimental requirements.
We remark that all strategies but Strategy 1 are characterized by having the initial vacuum modes initialized by active gates $S_{1}$ and $S_{2}$, followed by only passive gates. This approach has some advantages. First of all, by virtue of Euler’s decomposition in Eq.˜26, this approach is non-restrictive as it still enables us to explore the whole set of circuits that can be generated by the full set of gates. Second, we avoid circuits with multiple squeezers in sequence, which not only may be experimentally challenging but can also cause numerical errors due to a compounded large squeezing parameter, thereby invalidating the results.
@@ B.2 Results
We run the random search and reinforcement learning routine over a fixed number of episodes, where each episode terminates when the circuit depth reaches its maximum predefined value, $n_{\rm circuit}$. We explored different values of $n_{\rm circuit}$ to verify what violations can be achieved both in circuits with few gates and in circuits with large depth. The parameters of the photonic gates are optimized at the end of each episode such that the CHSH score is maximized, with a numerical procedure based on Nelder-Mead optimization. Subsequently, we fix the CHSH score and optimize the gates parameters to maximize the total heralding probability of the state.
In Table˜4 and Table˜5, we report a selection of photonic circuits found by random search and reinforcement learning, characterized by their CHSH score and heralding probability. For each search strategy, we report only the best circuit across both heralding mechanisms, omitting circuits based on a mechanism that is outperformed by the other. We focus on circuits with $N=4$ and $N=6$ modes, as our findings suggest that five modes offer no significant advantage over four.
In order to evaluate the various strategies without having to fine tune the RL parameters (e.g. update frequency, trajectory capacity, clip range, ADAM optimizer, and so on), we perform a random search, where the photonic operations are selected uniformly at random in each step, thereby saving us some time. In this evaluation, strategies 2 and 4 appear suboptimal (c.f. Table˜4), and we thus focus on the other strategies.
@@
Figure 5: The evolution of the CHSH score with the number of completed episodes, for an agent with PPO policy implementing three different strategies to build the optical circuits (c.f. Appendix˜B). In both (a) and (b), the increase of the CHSH score with the number of episodes shows that the agent learns from previous interactions with the environment to produce circuits with higher CHSH violations. In particular, the agent finds circuits with CHSH score equal to 2.068 (red dashed line) in (a) and 2.072 (red dashed line) in (b). Conversely, the CHSH score remains constantly below 2.0 (red line) in (c), signifying that the agent could not learn a policy leading to CHSH violations. Samples of the circuits found with the three strategies are reported in Table˜5.
@@
The circuits found by the agent in Table˜5 are representative of the learning process occurring with the PPO policy trained using the hyperparameters given in Table˜3. In fact, Table˜5 reports circuits found by the agent towards the end of the training process, where the CHSH score associated with the circuit becomes larger. The learning process can be visualized with the plots in Fig.˜5, which show the progression of the CHSH score over the training episodes. We observe that, for Strategy 1 in Fig.˜5(a) and Strategy 3 in Fig.˜5(b), the CHSH score improves over the training and reaches the maximum value in the later episodes. Conversely, it seems that the PPO agent struggles to learn circuit patterns leading to CHSH violations when adopting Strategy 5. This could be explained by the fact that, in Strategy 5, Alice’s and Bob’s modes are initialized in a product state that would result, if measured, in uncorrelated measurement outcomes ($\braket{A_{x}B_{y}}=0$), opposed to the correlations that arise when initializing the modes with a two-mode squeezer ($\braket{A_{x}B_{y}}>0$). This means that the starting value of the CHSH score, when the circuit is still empty, with Strategy 5 is closer to zero compared to Strategy 3, where it starts closer to $2$. Thus, it is easier for the agent in Strategy 3 to find circuit patterns leading to CHSH violations than when Strategy 5 is adopted. Still, we observe that the PPO agent in Strategy 5 can find circuits leading to a CHSH score of 2.0, i.e., a policy that generate (trivial) classical correlations between Alice and Bob is learned.
Parameter Value
Optimizer Adam ($1\times 10^{-4}$)
Discount factor $\gamma$ 0.99
GAE parameter $\lambda$ 0.95
Clip range 0.1
Max gradient norm 0.5
Number of epochs 25
Number of microbatches 4
Actor loss weight 1.0
Critic loss weight 0.5
Entropy loss weight 0.001
Update frequency 64
Trajectory Capacity 256
Table 3: Hyperparameters used to train PPO agents as defined in [60].
@@
Finally, training the agent both with the PPO policy and the random search allows us to benchmark the usefulness of a learning agent for circuit generation. By comparing the circuits found in Tables 4 and 5, there is no apparent advantage in using a trained agent to find good candidate circuits, compared to a random search, in this scenario. We remark, however, that this conclusion holds in the context of our RL trainings, which are limited by the computational power of a personal computer and by a shallow neural network with two hidden layers, but we expect a larger neural network to perform better than the one used here. Furthermore, we expect the performance of an unstructured search to be less appealing when trying to construct circuits in larger scenarios, e.g. involving more than two parties.
@@
Table 4: Circuits found by the agent with random search. “S.” indicates the strategy adopted, $n_{\text{circuit}}$ is the number of Gaussian operations, where we indicate in brackets the number of active Gaussian operations (squeezers). The notation is the one adopted in Table˜1 and the gates are applied sequentially in the same order as they appear in the list. We omit the $\hat{}$ symbol in the gates list for ease of notation. The heralding of the modes $3,4,\dots,N$ is “cl.” when the state is heralded on a click in a threshold detector, and “s.p.” when attained by simulating a single photon projection; we also report the total heralding probability of the state under “Herald.”.
S. NN $n_{\text{circuit}}$ Gates Herald. CHSH
@@
S. NN $n_{\text{circuit}}$ Gates Herald. CHSH
@@
Table 5: Circuits found using reinforcement learning with the PPO policy. The same notation of Table˜4 applies.
S. NN $n_{\text{circuit}}$ Gates Herald. CHSH